OpenStudy (anonymous):

Find f(3) and f'(3), assuming that the tangent line to y = f(x) at a = 3 has equation y = 6x + 8

OpenStudy (dumbcow):

the slope of the tangent line equals the derivative at the point x=3 -> f'(3) = 6 at x=3 both f(x) and the tangent line share the same point y = 6(3) +8 = 26 -> f(3) = 26

OpenStudy (anonymous):

so f ' (3) would be what? 1?

OpenStudy (campbell_st):

f(x) = 6x + 8 then f ' (x) = 6 so the gradient is a constant value.... m = 6.... find f(3) = 6 x 3 + 8 f(3) 26 f(-3) = 6 x (-3) + 8 f(-3) = -10 so you need to find 2 lines m = 6 point (3, 26) and m = 6 point (-3, -10)